# Multi-Objective Agentic Rewrites for Unstructured Data Processing

### v1

@@ 4.2. Selection
After initialization, the search tree contains $|V_{t}|=|M|+2|\hat{\mathcal{F}}_{|M|}|$ pipelines (counting toward the budget $B$). Now, in each iteration, the selection phase identifies which pipeline $P^{\star}$ to rewrite—equivalently, which path in the search tree to extend by one more rewrite. The selector must balance exploitation of high-performing paths with exploration of under-tested ones. MOAR achieves this balance by assigning each pipeline a multi-armed bandit-style utility score (auer2002finite).
@@
One widely used strategy for multi-armed bandits is the Upper Confidence Bound (UCB) algorithm (auer2002finite). However, we cannot apply UCB directly because the pipelines are not independent; each is derived from rewrites of its ancestors. Instead, we use UCT (kocsis2006bandit), which extends UCB to a tree-structured search space. Following UCT, we define the utility score as:
@@
Hierarchical Traversal with Progressive Widening. Because utility scores depend on $n_{t}(\mathrm{parent}(P))$, they are only comparable among siblings. We therefore traverse the tree top-down to select $P^{\star}$: starting at $P_{0}$, we repeatedly select the child with the highest utility until reaching a node $P^{\star}$ that can generate a new child. In standard UCT (kocsis2006bandit), when a node is selected, all possible rewrites (i.e., actions) are applied before the next iteration of selection. This is feasible for small action spaces (e.g., board games), but problematic here since a single pipeline could spawn hundreds of children (30+ directives $\times$ multiple operators $\times$ multiple instantiations). To limit branching, we use progressive widening, a technique from Monte Carlo Tree Search for large action spaces (chaslot2008progressive). The idea is to cap the number of children based on visit count, allowing the action space to be explored gradually. We accordingly set the maximum number of children for node $P$ to $W(n_{t}(P))=\max(2,\,1+\sqrt{n_{t}(P)})$, where the $\sqrt{n_{t}(P)}$ growth rate is typical (chaslot2008progressive). For example, a node with four visits may have at most three children; only after nine visits may it produce a fourth. This sublinear growth forces exploration of deeper paths before generating more rewrites from any single node. Algorithm 2 in App. C specifies the complete procedure.
@@ 4.3. Rewriting and Evaluation
Given the selected pipeline $P^{\star}$, the rewriting phase generates a new child pipeline $P^{\prime}$ by applying a directive. In DocETL-V1 (shankar2025docetl), optimization enumerates all applicable directives for each operator (or operator prefix) in the pipeline, then invokes an LLM agent to instantiate each one. The exhaustive approach is infeasible for MOAR, because the rewrite space is too large. Instead, MOAR delegates the entire rewriting decision to an LLM agent, which can reason about pipeline semantics to choose which directive to apply, which operators within the pipeline to target, and how to instantiate the directive. Algorithm 3 in App. C details the rewriting and evaluation procedure. Unlike selection, rewriting and evaluation are fully parallelized—multiple workers can simultaneously rewrite different selected pipelines and execute them on $D_{o}$. Our implementation uses 3 workers (capped by LLM API rate limits). We now describe how directives are encoded, how the agent chooses and instantiates them, and how resulting pipeline(s) are evaluated.
@@ 4.3.1. How Directives Are Encoded
  • Instantiation schema. A Pydantic model (pydantic) specifying parameters needed to apply the directive (e.g., clarify_instructions requires a clarified_prompt), with optional validators to enforce constraints (e.g., that the new prompt preserves all input variables).
@@ 4.3.2. Choosing and Instantiating Directives
Our agent-based rewriting proceeds in two steps. First, the agent chooses which directive to apply and which operators to target, seeing only directive names, descriptions, and use case guidance. Second, the agent loads the full instantiation schema and example application to generate concrete parameters. This progressive disclosure—a technique in user interface design for reducing cognitive load by revealing information gradually (CarrollRosson1987ProgressiveDisclosure)—avoids overwhelming the agent with details. [^3] Claude Code uses a similar approach for presenting documentation to agents (ZhangLazukaMurag2025).
@@ 5  Evaluation
We evaluate MOAR across a diverse set of workloads spanning legal, biomedical, government, consumer, and corporate domains. MOAR discovers the highest-accuracy pipeline on every workload. On average, MOAR achieves 27% higher accuracy than the next-best optimizer, ABACUS (russo2025abacus), while matching its best accuracy at only 54.5% of its cost. We first describe our setup (Sec. 5.1). We then present the accuracy improvements and cost savings achieved by MOAR (Sec. 5.2). Finally, we examine the characteristics of high-accuracy pipelines discovered by MOAR (Sec. 5.3). Our experiment artifacts are released at this Google Drive link.
@@ 5.1. Setup
We run MOAR with a budget of 40 pipeline evaluations per workload, using gpt-5 as the agent for instantiating rewrite directives. MOAR selects models from a pool of 11 LLMs, including gpt-4o-mini, gpt-4o, 3 gpt-4.1 variants, 3 gpt-5 variants, and 3 gemini-2.5 variants. All models are available to all optimizers. Unless otherwise noted, the user-specified or initial pipelines prior to optimization use gpt-4o-mini, as in prior work (russo2025abacus; shankar2025docetl; patel2025semantic).
@@ 5.1.1. Baselines
We compare MOAR against 4 baseline systems, including the original DocETL-V1 optimizer and a naive “agentic” baseline. We compare against open-source systems that support a semantic map operator. For each baseline, we express the pipeline using the baseline’s interface, minimally modifying the DocETL operators so the pipeline can be parsed by the baseline.
  • DocETL-V1 (shankar2025docetl). We run the original DocETL optimizer, which optimizes only for accuracy. DocETL-V1 returns a single plan.
  • Simple agent. We test whether an LLM agent can discover effective rewrites for DocETL, without explicit directives or structured search provided by MOAR. We provide a gpt-5 agent with three tools: (i) reading sample documents, (ii) reading DocETL documentation or the original DocETL paper (shankar2025docetl), and (iii) executing any pipeline on samples and observing its accuracy and cost. The agent proposes rewrites until its context window is exhausted ($\approx$400k tokens) or it calls a “done” tool to indicate completion. Then, of all pipelines generated by the agent, we retain the Pareto frontier.
  • LOTUS (patel2025semantic). We express each pipeline using LOTUS’s semantic operators. LOTUS does not support structured output schemas, so we minimally edit prompts to request JSON-formatted responses and write custom Python code to parse LOTUS outputs. LOTUS performs cost reduction for filters, joins, and group-bys by swapping in cheaper models (i.e., gpt-5-nano). Like DocETL-V1, LOTUS always returns a single optimized plan.
  • Palimpzest/ABACUS (liu2024declarative; russo2025abacus). We express each pipeline using Palimpzest (PZ)’s (liu2024declarative) operators and use its ABACUS optimizer (russo2025abacus). ABACUS does not directly return a Pareto frontier; instead, it allows users to specify a cost budget and returns the maximum-accuracy plan within that budget. Following discussion with the authors, we construct a Pareto frontier by running PZ with: (i) no budget constraint (to obtain the highest-accuracy plan), and (ii) budgets set to 50%, 25%, and 10% of the unconstrained plan’s cost. As of writing this paper (October 15, 2025), PZ does not support LLM-powered reduce operators, so we omit it from workloads requiring the DocETL reduce operator.
@@ 5.1.2. Workloads
We evaluate across six workloads. We take text processing workloads identified in prior published work (shankar2025docetl; patel2025semantic; russo2025abacus), and introduce two new workloads from medical analysis and enterprise sustainability domains, inspired by real DocETL users. The Sustainability workload is inspired by the Scottish Climate Intelligence Service, who uses DocETL to identify common sustainability initiatives across organizations and regional authorities. The MEDEC workload is inspired by AnkiHub, who builds study tools for medical students (e.g., personalized quizzes and automated error detection in their reasoning). For privacy, we use public datasets with similar task structures rather than proprietary user data.
@@
For each workload, we describe the task, initial pipeline, and metric. The initial pipeline for each workload uses a minimal number of semantic operators, representing what a user would naturally write when first encountering the problem (shankar2025steering). Each workload uses a dataset $D$ sampled from larger source datasets, split into $D_{o}$ (40 documents per workload, for optimization) and a held-out test set $D_{T}=D\setminus D_{o}$ (100 documents per workload). Optimizers search over candidate pipelines using $D_{o}$, and all reported results reflect execution on the held-out test set $D_{T}$.
@@
  • CUAD (Legal Analysis) (hendrycks2021cuad). This dataset contains 510 legal contracts (each averaging 7,727 words) annotated with 41 clause categories, used as a benchmark in prior work (shankar2025docetl; russo2025abacus). The task is to extract text spans for each clause type present in a contract. The initial pipeline consists of a single map that prompts for all clause types at once and outputs a list of {clause_type, text_span } objects. For PZ, we use the pipeline provided by the ABACUS authors, doubling the number of samples (as mentioned in Sec. 5.1.1) to aid PZ’s exploration. The metric is F1 score, counting an extraction as correct if the clause type matches and the span has Jaccard similarity >0.15>0.15 with the ground truth.
@@
  • Game Reviews. This dataset, from prior work (shankar2025docetl), involves documents from the Steam reviews dataset (sobkowicz2016steam), where each document contains 300 reviews for a single game (averaging 97,696 words). The task is to identify ten positive and ten negative reviews per game, presented in chronological order. The initial pipeline, as in  (shankar2025docetl), is a single map over each document that attempts to extract and order the required reviews directly. The metric is an average of hallucination rate (fraction of extracted reviews not present in the source text), sentiment accuracy (agreement with ground-truth ratings for non-hallucinated reviews), and Kendall’s $\tau$ between predicted and correct orderings. Since documents exceed context windows of most LLMs, we use only gpt-4.1 and gemini series models, with gpt-4.1-mini as default. We note that PZ crashed during optimization with a 200-sample budget ($\sim$4 hours in), so we reduced PZ’s budget to 50 samples.
@@
  • BlackVault (Declassified Articles). This dataset from (shankar2025docetl) contains 733 articles (each averaging 7,351 words) describing international paranormal events. The task is to classify each document’s event type and aggregate distinct location mentions across documents of the same type. The initial pipeline, as in (shankar2025docetl), has two operators: a map that extracts an event type per article (e.g., UFO sighting), followed by a reduce that aggregates locations by type. The metric is average recall of distinct locations per event type, normalized by the maximum recall achieved across all methods in Sec. 5.1.1 (to get a score between 0 and 1).
@@
  • Biodex (Biomedical Classification). This benchmark (d2023biodex) evaluates biomedical paper classification (each averaging 71,151 words). Each paper must be linked to the adverse drug reactions it discusses, drawn from over 24,000. Prior work expressed this task using different pipelines (code provided by the respective paper authors). LOTUS-r&r (retrieve-and-rerank) implements a join (to find all reactions for each document) followed by a semantic aggregation (to group by document and rerank reactions with an LLM) (patel2025semantic). PZ-r&r implements a map, retrieve, and map pipeline (russo2025abacus). However, now that LLMs support sufficiently long context windows (all models in our pool exceed 128k tokens), we implement the simplest initial pipeline for MOAR: a single map operation where the prompt contains the full list of reactions and the output is a ranked list of reactions relevant to the document. For fair comparison, we also express this single-map initial pipeline in LOTUS and PZ, called LOTUS-d and PZ-d (direct). We report results for both pipeline formulations (d and r&r) for LOTUS and PZ, taking the union of plans found for their Pareto frontiers. The metric is rank-precision@5 (RP@5), measuring how often the correct reactions appear among the top-5 predictions.
@@
  • MEDEC (Medical Error Detection and Correction); new. The MEDEC dataset (abacha2024medec) contains 3,848 clinical notes (each averaging 147 words) with labeled medical errors across five categories. The task is to detect whether an error is present in each note, identify the sentence containing the error, and generate a corrected version. The initial pipeline is a single map that produces three outputs: error_flag, error_sentence, and corrected_sentence. The metric is an average of the error-detection F1 scores and the corrected sentence Jaccard similarities with the reference corrections.
@@
  • Sustainability (Corporate ESG Reports); new. This dataset (datanneed2024sustainability) contains 5,436 enterprise reports (each averaging 38,668 words) spanning annual reports, sustainability reports, financial reports, and others. The task is to (i) filter to retain only sustainability reports, (ii) classify each company’s economic sector (e.g., health, real estate), and (iii) for each sector, produce a summary listing each company and its key sustainability initiatives (e.g., for the technology sector: “Apple: carbon neutrality by 2030; Microsoft: 100% renewable energy; Google: water replenishment programs”). The initial pipeline applies a filter to select sustainability reports, a map to classify the economic sector, and a reduce to group reports by sector and generate a summary per sector containing all companies and their initiatives. The metric is the average of sector classification accuracy (the fraction of reports assigned to the correct sector) and company name accuracy (the fraction of company names in the per-sector summaries that match ground-truth companies for that sector).

### v3

@@ 4.2. Selection
After initialization, the search tree contains $|V_{t}|=|M|+2|\hat{\mathcal{F}}_{|M|}|$ pipelines (counting toward the budget $B$). Now, in each iteration, the selection phase identifies which pipeline $P^{\star}$ to rewrite—equivalently, which path in the search tree to extend by one more rewrite. The selector must balance exploitation of high-performing paths with exploration of under-tested ones. MOAR achieves this balance by assigning each pipeline a multi-armed bandit-style utility score (Auer et al., 2002).
@@
One widely used strategy for multi-armed bandits is the Upper Confidence Bound (UCB) algorithm (Auer et al., 2002). However, we cannot apply UCB directly because the pipelines are not independent; each is derived from rewrites of its ancestors. Instead, we use UCT (Kocsis and Szepesvári, 2006), which extends UCB to a tree-structured search space. Following UCT, we define the utility score as:
@@
Hierarchical Traversal with Progressive Widening. Because utility scores depend on $n_{t}(\mathrm{parent}(P))$, they are only comparable among siblings. We therefore traverse the tree top-down to select $P^{\star}$: starting at $P_{0}$, we repeatedly select the child with the highest utility until reaching a node $P^{\star}$ that can generate a new child. In standard UCT (Kocsis and Szepesvári, 2006), when a node is selected, all possible rewrites (i.e., actions) are applied before the next iteration of selection. This is feasible for small action spaces (e.g., board games), but problematic here since a single pipeline could spawn hundreds of children (30+ directives $\times$ multiple operators $\times$ multiple instantiations). To limit branching, we use progressive widening, a technique from Monte Carlo Tree Search for large action spaces (Chaslot et al., 2008). The idea is to cap the number of children based on visit count, allowing the action space to be explored gradually. We accordingly set the maximum number of children for node $P$ to $W(n_{t}(P))=\max(2,\,1+\sqrt{n_{t}(P)})$, where the $\sqrt{n_{t}(P)}$ growth rate is typical (Chaslot et al., 2008). For example, a node with four visits may have at most three children; only after nine visits may it produce a fourth. This sublinear growth forces exploration of deeper paths before generating more rewrites from any single node. Algorithm 2 in App. C specifies the complete procedure.
@@ 4.3. Rewriting and Evaluation
Given the selected pipeline $P^{\star}$, the rewriting phase generates a new child pipeline $P^{\prime}$ by applying a directive. In DocETL-V1 (Shankar et al., 2025a), optimization enumerates all applicable directives for each operator (or operator prefix) in the pipeline, then invokes an LLM agent to instantiate each one. The exhaustive approach is infeasible for MOAR, because the rewrite space is too large. Instead, MOAR delegates the entire rewriting decision to an LLM agent, which can reason about pipeline semantics to choose which directive to apply, which operators within the pipeline to target, and how to instantiate the directive. Algorithm 3 in App. C details the rewriting and evaluation procedure. Unlike selection, rewriting and evaluation are fully parallelized—multiple workers can simultaneously rewrite different selected pipelines and execute them on $D_{o}$. Our implementation uses 3 workers (capped by LLM API rate limits). We now describe how directives are encoded, how the agent chooses and instantiates them, and how resulting pipeline(s) are evaluated.
@@ 4.3.1. How Directives Are Encoded
  • Instantiation schema. A Pydantic model (Colvin and the Pydantic contributors, 2025) specifying parameters needed to apply the directive (e.g., clarify_instructions requires a clarified_prompt), with optional validators to enforce constraints (e.g., that the new prompt preserves all input variables).
@@ 4.3.2. Choosing and Instantiating Directives
Our agent-based rewriting proceeds in two steps. First, the agent chooses which directive to apply and which operators to target, seeing only directive names, descriptions, and use case guidance. Second, the agent loads the full instantiation schema and example application to generate concrete parameters. This progressive disclosure—a technique in user interface design for reducing cognitive load by revealing information gradually (Carroll and Rosson, 1987)—avoids overwhelming the agent with details. [^3] Claude Code uses a similar approach for presenting documentation to agents (Zhang et al., 2025).
@@ 5  Evaluation
We evaluate MOAR across a diverse set of workloads spanning legal, biomedical, government, consumer, and corporate domains. MOAR discovers the highest-accuracy pipeline on every workload. On average, MOAR achieves 27% higher accuracy than the next-best optimizer, ABACUS (Russo et al., 2025), while matching its best accuracy at only 54.5% of its cost. We first describe our setup (Sec. 5.1). We then present the accuracy improvements and cost savings achieved by MOAR (Sec. 5.2). Finally, we examine the characteristics of high-accuracy pipelines discovered by MOAR (Sec. 5.3). Our experiment artifacts are released at this Google Drive link.
@@ 5.1. Setup
We run MOAR with a budget of 40 pipeline evaluations per workload, using gpt-5 as the agent for instantiating rewrite directives. MOAR selects models from a pool of 11 LLMs, including gpt-4o-mini, gpt-4o, 3 gpt-4.1 variants, 3 gpt-5 variants, and 3 gemini-2.5 variants. All models are available to all optimizers. Unless otherwise noted, the user-specified or initial pipelines prior to optimization use gpt-4o-mini, as in prior work (Russo et al., 2025; Shankar et al., 2025a; Patel et al., 2025).
@@ 5.1.1. Baselines
We compare MOAR against 4 baseline systems, including the original DocETL-V1 optimizer and a naive “agentic” baseline. We compare against open-source systems that support a semantic map operator. For each baseline system, we express the pipeline using the baseline’s interface, minimally modifying the DocETL operators so the pipeline can be parsed by the baseline system. All baselines except DocETL-V1 support selecting from multiple models; DocETL-V1 uses only gpt-4o-mini.
  • DocETL-V1 (Shankar et al., 2025a). We run the original DocETL optimizer, which optimizes only for accuracy and returns a single plan. Unlike all other baselines, DocETL-V1 does not access ground-truth labels; it relies solely on an LLM judge to select plans.
  • Simple agent. We test whether an LLM agent can discover effective DocETL pipelines, without the structured search provided by MOAR. We provide a gpt-5 agent with three tools: (i) reading sample documents, (ii) reading the original DocETL paper (Shankar et al., 2025a) or DocETL documentation, which contains many example pipelines and rewrite descriptions, and (iii) executing any pipeline on samples and observing its accuracy and cost. The agent proposes rewrites until its context window is exhausted ($\approx$400k tokens) or it calls a “done” tool to indicate completion. Then, of all pipelines generated by the agent, we retain the Pareto frontier.
  • LOTUS (Patel et al., 2025). We express each pipeline using LOTUS’s semantic operators. LOTUS does not support structured output schemas, so we minimally edit prompts to request JSON-formatted responses and write custom Python code to parse LOTUS outputs. LOTUS performs cost reduction for filters, joins, and group-bys by swapping in cheaper models (i.e., gpt-5-nano). Like DocETL-V1, LOTUS always returns a single optimized plan.
  • Palimpzest/ABACUS (Liu et al., 2025a; Russo et al., 2025). We express each pipeline using Palimpzest (PZ)’s (Liu et al., 2025a) operators and use its ABACUS optimizer (Russo et al., 2025). ABACUS does not directly return a Pareto frontier; instead, it allows users to specify a cost budget and returns the maximum-accuracy plan within that budget. Following discussion with the authors, we construct a Pareto frontier by running PZ with: (i) no budget constraint (to obtain the highest-accuracy plan), and (ii) budgets set to 50%, 25%, and 10% of the unconstrained plan’s cost. As of writing this paper (October 15, 2025), PZ does not support LLM-powered reduce operators, so we omit it from workloads requiring the DocETL reduce operator.
@@ 5.1.2. Workloads
We evaluate across six workloads. We take text processing workloads identified in prior published work (Shankar et al., 2025a; Patel et al., 2025; Russo et al., 2025), and introduce two new workloads from medical analysis and enterprise sustainability domains, inspired by real DocETL users. The Sustainability workload is inspired by the Scottish Climate Intelligence Service, who uses DocETL to identify common sustainability initiatives across organizations and regional authorities. The MEDEC workload is inspired by AnkiHub, who builds study tools for medical students (e.g., personalized quizzes and automated error detection in their reasoning). For privacy, we use public datasets with similar task structures rather than proprietary user data.
@@
For each workload, we describe the task, initial pipeline, and metric. The initial pipeline for each workload uses a minimal number of semantic operators, representing what a user would naturally write when first encountering the problem (Shankar et al., 2025b). Each workload uses a dataset $D$ sampled from larger source datasets, split into $D_{o}$ (40 documents per workload, for optimization) and a held-out test set $D_{T}=D\setminus D_{o}$ (100 documents per workload). Optimizers search over candidate pipelines using $D_{o}$, and all reported results reflect execution on the held-out test set $D_{T}$.
@@
  • CUAD (Legal Analysis) (Hendrycks et al., 2021). This dataset contains 510 legal contracts (each averaging 7,727 words) annotated with 41 clause categories, used as a benchmark in prior work (Shankar et al., 2025a; Russo et al., 2025). The task is to extract text spans for each clause type present in a contract. The initial pipeline consists of a single map that prompts for all clause types at once and outputs a list of {clause_type, text_span } objects. For PZ, we use the pipeline provided by the ABACUS authors, doubling the number of samples (as mentioned in Sec. 5.1.1) to aid PZ’s exploration. The metric is F1 score, counting an extraction as correct if the clause type matches and the span has Jaccard similarity >0.15>0.15 with the ground truth.
@@
  • Game Reviews. This dataset, from prior work (Shankar et al., 2025a), involves documents from the Steam reviews dataset (Sobkowicz and Stokowiec, 2016), where each document contains 300 reviews for a single game (averaging 97,696 words). The task is to identify ten positive and ten negative reviews per game, presented in chronological order. The initial pipeline, as in  (Shankar et al., 2025a), is a single map over each document that attempts to extract and order the required reviews directly. The metric is an average of hallucination rate (fraction of extracted reviews not present in the source text), sentiment accuracy (agreement with ground-truth ratings for non-hallucinated reviews), and Kendall’s $\tau$ between predicted and correct orderings. Since documents exceed context windows of most LLMs, we use only gpt-4.1 and gemini series models, with gpt-4.1-mini as default. We note that PZ crashed during optimization with a 200-sample budget ($\sim$4 hours in), so we reduced PZ’s budget to 50 samples.
@@
  • BlackVault (Declassified Articles). This dataset from (Shankar et al., 2025a) contains 733 articles (each averaging 7,351 words) describing international paranormal events. The task is to classify each document’s event type and aggregate distinct location mentions across documents of the same type. The initial pipeline, as in (Shankar et al., 2025a), has two operators: a map that extracts an event type per article (e.g., UFO sighting), followed by a reduce that aggregates locations by type. The metric is average recall of distinct locations per event type, normalized by the maximum recall achieved across all methods in Sec. 5.1.1 (to get a score between 0 and 1).
@@
  • Biodex (Biomedical Classification). This benchmark (D’Oosterlinck et al., 2023) evaluates biomedical paper classification (each averaging 71,151 words). Each paper must be linked to the adverse drug reactions it discusses, drawn from over 24,000. Prior work expressed this task using different pipelines (code provided by the respective paper authors). LOTUS-r&r (retrieve-and-rerank) implements a join (to find all reactions for each document) followed by a semantic aggregation (to group by document and rerank reactions with an LLM) (Patel et al., 2025). PZ-r&r implements a map, retrieve, and map pipeline (Russo et al., 2025). However, now that LLMs support sufficiently long context windows (all models in our pool exceed 128k tokens), we implement the simplest initial pipeline for MOAR: a single map operation where the prompt contains the full list of reactions and the output is a ranked list of reactions relevant to the document. For fair comparison, we also express this single-map initial pipeline in LOTUS and PZ, called LOTUS-d and PZ-d (direct). We report results for both pipeline formulations (d and r&r) for LOTUS and PZ, taking the union of plans found for their Pareto frontiers. The metric is rank-precision@5 (RP@5), measuring how often the correct reactions appear among the top-5 predictions.
@@
  • MEDEC (Medical Error Detection and Correction); new. The MEDEC dataset (Abacha et al., 2024) contains 3,848 clinical notes (each averaging 147 words) with labeled medical errors across five categories. The task is to detect whether an error is present in each note, identify the sentence containing the error, and generate a corrected version. The initial pipeline is a single map that produces three outputs: error_flag, error_sentence, and corrected_sentence. The metric is an average of the error-detection F1 scores and the corrected sentence Jaccard similarities with the reference corrections.
@@
  • Sustainability (Corporate ESG Reports); new. This dataset (DataNeed, 2024) contains 5,436 enterprise reports (each averaging 38,668 words) spanning annual reports, sustainability reports, financial reports, and others. The task is to (i) filter to retain only sustainability reports, (ii) classify each company’s economic sector (e.g., health, real estate), and (iii) for each sector, produce a summary listing each company and its key sustainability initiatives (e.g., for the technology sector: “Apple: carbon neutrality by 2030; Microsoft: 100% renewable energy; Google: water replenishment programs”). The initial pipeline applies a filter to select sustainability reports, a map to classify the economic sector, and a reduce to group reports by sector and generate a summary per sector containing all companies and their initiatives. The metric is the average of sector classification accuracy (the fraction of reports assigned to the correct sector) and company name accuracy (the fraction of company names in the per-sector summaries that match ground-truth companies for that sector).